Analytic Geometry
5. Determine the equation of the line in standard form:
a) with a slope of -3/2 and passing through the point (5, -1).
Thank you very much!!!

Answer by checkley75(3666)

Y=mX+b where m=slope & b=the y intercept.
-1=(-3/2)5+b
-1=-15/2+b
-1+15/2=b
-1+7.5=b
6.5=b ans. for the y intercept.
y=-3x/2+6.5